OpenStudy (anonymous):

Explain, in complete sentences, the relationship between the equations of perpendicular lines.

OpenStudy (nurali):

The relationship between perpendicular lines are based off of their slope. The slope of one line is the opposite reciprocal of the other. [ Written mathematically, the lines y=m*x +b and y =(-1/m)*x +c are perpendicular lines (note the y-intercepts do not need to be equal or even related to each other).

OpenStudy (anonymous):

Perpendicular lines intersect each other at a 90 degree angle. A perfect example of perpendicular lines are the x and y axis of a coordinate plane system. Hope that helps!! Remember if you are taking Geometry right now that there are different theorems you might have to know for proof problems.

OpenStudy (anonymous):

They have both touched it fairly well, but the way you get a perpendicular line, which nurali briefly said, is that you take the negative reciprocal of the slope. That is to say, you look at the slope from the first equation, ie y=mx+b where "m" is the slope. You take its negative reciprocal which would be -(1/m). That is the slope for your new line. From there, just plug it into the point slope form: y-y1=m(x-x1) where y1 is a y coordinate of the new line, and x1 is an x coordinate on the new line.